The length of a rectange is (x+4)cm.
The width is (x-3)cm.
The area of the rectangle is 78cm^2.
How do I write this in terms of x?

Answer by richwmiller(17219)

(x+4)*(x-3)=78
x^2+x-90 = 0
x=-10 x=9
we discard x=-10
9+4=13
9-3-6